# Holon metal, charge-density-wave and chiral superconductor from doping fractional Chern insulator and SU(3)1 chiral spin liquid

### v1

@@ Abstract
Recent experiment observed a superconductor proximate to the $\nu_{h}=-\frac{2}{3}=-1+\frac{1}{3}$ fractional quantum anomalous Hall (FQAH) insulator in twisted MoTe2. One critical question is whether the normal state is a Fermi liquid with large Fermi surface, or a strongly correlated metallic state with small carrier density. In this work we develop a theory of possible phases from doping the $\nu=-\frac{2}{3}$ fractional Chern insulator (FCI). We also point out that the problem is dual to doping a SU(3)1 chiral spin liquid(CSL) if we assume that the spin gap is finite. For both problems, a natural phase upon doping is a holon metal with three small Fermi pockets. The pocket is formed by a spinless charge $-e$ holon in the doped CSL case and a charge $-e/3$ fractionalized hole in the doped FCI case. The holon metal is likely unstable to pairing due to gauge field fluctuations, leading to a charge density wave (CDW) metal, but may be stabilized by magnetic field and shows an unusual quantum oscillation period. Two different chiral superconductors are possible, emerging from the CDW metal phase or directly from the holon metal. We note that a superconductor directly from anyon gas seems unlikely if the cheapest anyon is the elementary one with $e/3$ charge.
@@ I Introduction
To further demonstrate the duality, we derive the same low energy effective field theory for both problems. In the CSL side we use the standard slave boson theory and find the slave boson has three minima in its dispersion, as required by the projective translation symmetry. At low energy each minima provides a bosonic field $\varphi_{I}$, $I=1,2,3$, which feels an emergent internal flux and is at effective filling $\nu_{m}=-1$. For the doped SU(2)1 CSL, we only have two flavors and they can form a bosonic integer quantum Hall state, leading to a superconductor for the electron[46]. In contrast, now the only translation invariant state is to put each flavor $\varphi_{I}$ in a bosonic composite Fermi liquid (CFL) or Pfaffian state. Let us focus on the simplest CFL scenario. The final phase in terms of electron turns out to be a holon metal[48], which has three small Fermi pockets formed by spinless charge $-e$ holon. In the dual side of doped FCI, we can formulate the same effective theory, where $\varphi_{I}$ is now the vortex of the composite boson in the standard theory of Laughlin state. In the end if we put each $\varphi_{I}$ in a bosonic CFL state, again we get an exotic metal with three small pocket, but now formed by charge $-e/3$ fractionalized hole. For simplicity, we still dub the phase as holon metal. A similar phase was proposed in Ref. [40] for doped FCI and was called ‘secondary CFL’. However, the property of the phase is quite different from a CFL: (1) the phase is compressible and well defined in a range of density, while the CFL phase is restricted to a specific doping; (2) The phase has a finite drude weight and the conductivity $\sigma_{xx}$ could be arbitrarily large in the clean limit. In contrast the CFL phase has vanishing $\sigma_{xx}$ in the clean limit. For this reason we prefer the name holon metal. Our holon metal is a simpler version of the state in Ref. [40]. Intuitively this exotic metal may be conveniently constructed in this way: we fractionalize an electron into three fractionalized fermions with charge $e/3$, and then put each fermion in a $C=-1$ Chern insulator plus a small hole pocket.
@@ II.1 Doped SU(2)1 CSL
where $ada$ is the abbreviation of the Chern-Simons term $\epsilon_{\mu\nu\sigma}a_{\mu}\partial_{\nu}a_{\sigma}$. Here $\epsilon$ is the anti-symmetric tensor.
@@
Next we dope the system so the electron is at total filling $n=1-x$. We have $n_{b}=n_{f}=1-x$ on average. Now naively the spinon $f$ can not stay in a gapped Chern insulator phase and the spin gap is closed. In this simple case we expect that the slave boson is condensed and we just have a simple Fermi liquid coexisting with charge-density-wave (CDW) with doubled unit cell[46]. However, there is actually a possibility that the spin gap remains finite and the low energy physics is dominated by finite density of spinless holons $b$. We will focus on this scenario with a finite spin gap in this work. Because the spin is gapped, the low energy degree of freedom is actually dominated by the physical charge 2e Cooper pair $\Delta=c_{\uparrow}c_{\downarrow}$.
To keep the spinon $f_{\sigma}$ in a gapped Chern insulator phase, we need to have a finite internal flux $\frac{\langle da\rangle}{2\pi}=-\frac{x}{2}$ per unit cell. Here $da$ indicates $\partial_{x}a_{y}-\partial_{y}a_{x}$. With the dynamically generated internal flux, $f_{\sigma}$ stays in the Chern insulator phase following the Streda’s formula. Then the low energy theory is entirely from the slave boson $b$. $b$ is in a Mott insulator at $n=1$. Now at filling $n=1-x$, we basically dope its holes $\varphi\sim b^{\dagger}$ with density $n_{\varphi}=x$. However, due to the projective translation symmetry, $b$ feels a $\pi$ flux per triangular unit cell and the magnetic translation symmetry guarantees two minima. If one of them is at momentum $\mathbf{k}_{1}=(0,0)$, the other one is at $\mathbf{k}_{2}=(\pi,0)$. We label the low energy operator of $\varphi$ around these two minima as $\varphi_{1}(t,\mathbf{r})$ and $\varphi_{2}(t,\mathbf{r})$. Then the physical translation symmetry acts as: $T_{x}:\varphi_{1}\rightarrow\varphi_{1},\varphi_{2}\rightarrow\varphi_{2}$, $T_{y}:\varphi_{1}\rightarrow\varphi_{2},\varphi_{2}\rightarrow\varphi_{1}$. Each of them then is at density $n=\frac{x}{2}$ per triangular unit cell.
@@
Now let us tune $\mu$ to dope the CSL so that $\langle n_{\varphi_{I}}\rangle=\frac{x}{2}$. From variation of $a_{0}$ we can find $\frac{\langle da\rangle}{2\pi}=-\frac{x}{2}$ as expected. Next we need to discuss the ground state with finite density of $\varphi_{I}$. Due to the effective magnetic flux, simple boson condensation is not an option. Instead, $\varphi$ can form a bosonic integer quantum Hall (bIQH) state with $C=-2$. In the end this provides an additional $\frac{2}{4\pi}ada$ contribution from the bIQH phase, which cancels the self Chern-Simons term in the above Lagrangian. In the end the low energy theory is:
@@ II.2 Doped bosonic Laughlin state
One may view $\varphi^{\dagger}$ as the creation operator of the anyon with charge $e/2$ in the standard effective theory of the Laughlin state. One may view $\varphi$ as the vortex of the composite boson. When the physical boson is at filling $n=\frac{1}{2}-x$, $\varphi$ is at density $n_{\varphi}=2x$. This is also in agreement with the doped CSL theory where the filling of the electron is $n=1-2x$ so the Cooper pair density is $\frac{1}{2}-x$.
The next step is exactly the same as the doped CSL case. $\varphi$ feels an effective magnetic flux from $da$ and is at magnetic filling $\nu=-2$, so it can form a bosonic IQH phase with $C=-2$. In the end the final theory is just:
@@ III Holon metal from doped SU(3)1 CSL
To describe the CSL, we use the slave boson construction $c_{i;a}=b_{i}f_{i;a}$ with $a=1,2,3$ as the flavor index. $f_{a}$ is put into a $C=1$ insulator with a $\frac{2\pi}{3}$ flux per unit cell. Similarly, the slave boson $b$ will feel a $-2\pi/3$ flux per unit cell. Its dispersion will have three minima guaranteed by the projective translation symmetry. Next we dope the system so that the average density is $n_{b}=n_{f}=1-x$ per triangular unit cell. Similar to the SU(2) CSL case, $b$ and $f$ share an internal U(1) gauge field $a_{\mu}$. We assign the physical charge to the slave boson $b$. So in the end $b$ couples to $A_{c}-a$ and $f$ couples to $a$.
To keep $f$ in a $C=1$ Chern insulator, we need $\frac{\langle da\rangle}{2\pi}=-\frac{x}{3}$. Then the low energy theory is dominated by the slave boson $b$. Again $b$ feels an effective flux and its hole $\varphi\sim b^{\dagger}$ has three minima at momentum $\mathbf{k}_{1}=(0,0),\mathbf{k}_{2}=(\frac{2\pi}{3},0),\mathbf{k}_{3}=(\frac{4\pi}{3},0)$. We define the corresponding low energy field operator as $\varphi_{I}(t,x)$ with $I=1,2,3$. The physical translation symmetry acts as $T_{x}:\varphi_{I}\rightarrow\varphi_{I}$ and $T_{y}:\varphi_{1}\rightarrow\varphi_{2},\varphi_{2}\rightarrow\varphi_{3},\varphi_{3}\rightarrow\varphi_{1}$. The low-energy effective theory is:
@@
The translation symmetry acts as $T_{x}:\psi_{I}\rightarrow\psi_{I},\alpha_{\mu}\rightarrow\alpha_{\mu},\beta_{\mu}\rightarrow\beta_{\mu}$, $T_{y}:\psi_{1}\rightarrow\psi_{2},\psi_{2}\rightarrow\psi_{3},\psi_{3}\rightarrow\psi_{1},\alpha_{\mu}\rightarrow\beta_{\mu},\beta_{\mu}\rightarrow-\alpha_{\mu}-\beta_{\mu}$.
@@ III.1 Property of the holon metal
Here let us discuss the experimental property of the holon metal, with a focus on its transport signature. We hope to calculate the conductivity tensor $\sigma_{c}=\begin{pmatrix}\sigma_{xx}&\sigma_{xy}\\ -\sigma_{xy}&\sigma_{yy}\end{pmatrix}$. In the following, we assume that the translation symmetry is preserved. Thus $\psi_{1},\psi_{2},\psi_{3}$ has the same conductivity tensor $\sigma_{\psi_{1}}=\sigma_{\psi_{2}}=\sigma_{\psi_{3}}=\sigma_{\psi}$. It is easy to find the response to $A_{c}$ as:
@@
where we used the Coulomb gauge with $A_{0}=0$ and $\mathbf{q}\cdot\mathbf{A}(\omega,\mathbf{q})=0$. There is no crossing term between $A_{c}$ and $\alpha,\beta$.
$\Pi$ is the response from each $\psi_{I}$: $\Pi(\omega,\mathbf{q}=0)=\frac{e^{2}}{h}(-i\omega)\sigma_{\psi}=\frac{e^{2}}{h}(-i\omega)\begin{pmatrix}\sigma_{\psi;xx}&\sigma_{\psi;xy}\\ -\sigma_{\psi;-xy}&\sigma_{\psi;yy}\end{pmatrix}$. So in the end, we have a simple relationship: $\sigma_{c}=3\sigma_{\psi}$, provided that the translation symmetry is not broken. Therefore, the holon metal has a resistivity tensor which is very similar to a Fermi liquid. Because the time-reversal symmetry is broken, $\psi$ can still have a small non-universal Hall conductivity. But the final phase has a small Hall angle $\frac{\rho_{xy}}{\rho_{xx}}$ because $\sigma_{xx}$ is large in the clean system.
@@ III.2 Alternative derivation and wavefunction
To gain more insight into this exotic metallic phase, we also provide a different and perhaps simpler construction. In this phase, there are three pockets with Fermi surface size $A_{\mathrm{FS}}=\frac{x}{3}$ per Brillouin zone (BZ). We also find a gauge-invariant operator $C_{\mathrm{trion}}=\psi_{1}\psi_{2}\psi_{3}$ which is the physical trion operator. As discussed before, the physics is equivalent to doping a Laughlin state of the trion on top of a $C=3$ IQH phase. Now the trion is at density $n_{\mathrm{trion}}=\frac{1}{3}-\frac{x}{3}$, we simply do parton construction $C_{\mathrm{trion}}=f_{1}f_{2}f_{3}$ and then put each flavor $f_{I}$ in a $C=-1$ Chern insulator plus a small hole pocket formed by $\psi_{I}\sim f_{I}^{\dagger}$ with size $x/3$. In this construction, there is a SU(3) internal gauge field, but let us assume it’s higgsed down to $U(1)\times U(1)$ by some terms. In the end we only have the U(1) gauge field $\alpha_{\mu}$ and $\beta_{\mu}$. $\psi_{1}$ couples to $-A_{c}+\alpha_{\mu}$, $\psi_{2}$ couples to $-A_{c}+\beta_{\mu}$ and $\psi_{3}$ couples to $-A_{c}-\alpha_{\mu}-\beta_{\mu}$. The $C=-1$ Chern insulator gives a background Chern-Simons term. Eventually, we reproduce the theory in Eq. 14 once we also include the $\frac{3}{4\pi}A_{c}dA_{c}$ term from the stacked $C=3$ IQH phase.

### v2

@@ Abstract
Recent experiments have observed superconductivity proximate to the $\nu=-2/3$ fractional quantum anomalous Hall (FQAH) insulator in twisted MoTe2. A critical open question is whether the underlying normal state is a Fermi liquid with a large Fermi surface or a strongly correlated metal with low carrier density. In this work, we develop a theory of the phases emerging from doping the $\nu=-2/3$ fractional Chern insulator (FCI). We establish a duality between this problem and doping a gapped SU(3)1 chiral spin liquid (CSL). In both scenarios, one possible metallic state upon doping is a holon metal featuring three small Fermi pockets. These pockets are formed by spinless charge e-e holons in the CSL case and charge e/3-e/3 fractionalized holes in the FCI case. While the holon metal is susceptible to pairing instabilities driven by gauge fluctuations—leading to a charge density wave (CDW) metal—it may be stabilized by a magnetic field, where it would exhibit an anomalous quantum oscillation period. We identify two distinct chiral superconducting phases that can emerge either from the CDW metal or directly from the holon metal. Finally, we argue that superconductivity arising directly from an anyon gas is not likely if the lowest-energy anyon carries the elementary charge $e/3$.
@@ I Introduction
To further demonstrate the duality, we derive the same low energy effective field theory for both problems. In the CSL side we use the standard slave boson theory and find the slave boson has three minima in its dispersion, as required by the projective translation symmetry. At low energy each minima provides a bosonic field $\varphi_{I}$, $I=1,2,3$, which feels an emergent internal flux and is at effective filling $\nu_{m}=-1$. For the doped SU(2)1 CSL, we only have two flavors and they can form a bosonic integer quantum Hall state, leading to a superconductor for the electron[46]. In contrast, now the only translation invariant state is to put each flavor $\varphi_{I}$ in a bosonic composite Fermi liquid (CFL) or Pfaffian state. Let us focus on the simplest CFL scenario. The final phase in terms of electron turns out to be a holon metal[48], which has three small Fermi pockets formed by spinless charge e-e holon. In the dual side of doped FCI, we can formulate the same effective theory, where $\varphi_{I}$ is now the vortex of the composite boson in the standard theory of Laughlin state. In the end if we put each $\varphi_{I}$ in a bosonic CFL state, again we get an exotic metal with three small pocket, but now formed by charge e/3-e/3 fractionalized hole. For simplicity, we still dub the phase as holon metal. A similar phase was proposed in Ref. [40] for doped FCI and was called ‘secondary CFL’. However, the property of the phase is quite different from a CFL: (1) the phase is compressible and well defined in a range of density, while the CFL phase is restricted to a specific doping; (2) The phase has a finite drude weight and the conductivity $\sigma_{xx}$ could be arbitrarily large in the clean limit. In contrast the CFL phase has vanishing $\sigma_{xx}$ in the clean limit. For this reason we prefer the name holon metal. Our holon metal is a simpler version of the state in Ref. [40]. Intuitively this exotic metal may be conveniently constructed in this way: we fractionalize an electron into three fractionalized fermions with charge $e/3$, and then put each fermion in a C=1C=-1 Chern insulator plus a small hole pocket.
@@ II.1 Doped SU(2)1 CSL
where adaada is the abbreviation of the Chern-Simons term $\epsilon_{\mu\nu\sigma}a_{\mu}\partial_{\nu}a_{\sigma}$. Here $\epsilon$ is the anti-symmetric tensor.
@@
Next we dope the system so the electron is at total filling n=1xn=1-x. We have $n_{b}=n_{f}=1-x$ on average. Now naively the spinon $f$ can not stay in a gapped Chern insulator phase and the spin gap is closed. In this simple case we expect that the slave boson is condensed and we just have a simple Fermi liquid coexisting with charge-density-wave (CDW) with doubled unit cell[46]. However, there is actually a possibility that the spin gap remains finite and the low energy physics is dominated by finite density of spinless holons $b$. We will focus on this scenario with a finite spin gap in this work. Because the spin is gapped, the low energy degree of freedom is actually dominated by the physical charge 2e Cooper pair $\Delta=c_{\uparrow}c_{\downarrow}$.
To keep the spinon $f_{\sigma}$ in a gapped Chern insulator phase, we need to have a finite internal flux $\frac{\langle da\rangle}{2\pi}=-\frac{x}{2}$ per unit cell. Here $da$ indicates $\partial_{x}a_{y}-\partial_{y}a_{x}$. With the dynamically generated internal flux, $f_{\sigma}$ stays in the Chern insulator phase following the Streda’s formula. Then the low energy theory is entirely from the slave boson $b$. $b$ is in a Mott insulator at $n=1$. Now at filling n=1xn=1-x, we basically dope its holes $\varphi\sim b^{\dagger}$ with density $n_{\varphi}=x$. However, due to the projective translation symmetry, $b$ feels a $\pi$ flux per triangular unit cell and the magnetic translation symmetry guarantees two minima. If one of them is at momentum $\mathbf{k}_{1}=(0,0)$, the other one is at $\mathbf{k}_{2}=(\pi,0)$. We label the low energy operator of $\varphi$ around these two minima as $\varphi_{1}(t,\mathbf{r})$ and $\varphi_{2}(t,\mathbf{r})$. Then the physical translation symmetry acts as: $T_{x}:\varphi_{1}\rightarrow\varphi_{1},\varphi_{2}\rightarrow-\varphi_{2}$, $T_{y}:\varphi_{1}\rightarrow\varphi_{2},\varphi_{2}\rightarrow\varphi_{1}$. Each of them then is at density $n=\frac{x}{2}$ per triangular unit cell.
@@
Now let us tune $\mu$ to dope the CSL so that $\langle n_{\varphi_{I}}\rangle=\frac{x}{2}$. From variation of $a_{0}$ we can find $\frac{\langle da\rangle}{2\pi}=-\frac{x}{2}$ as expected. Next we need to discuss the ground state with finite density of $\varphi_{I}$. Due to the effective magnetic flux, simple boson condensation is not an option. Instead, $\varphi$ can form a bosonic integer quantum Hall (bIQH) state with C=2C=-2. In the end this provides an additional $\frac{2}{4\pi}ada$ contribution from the bIQH phase, which cancels the self Chern-Simons term in the above Lagrangian. In the end the low energy theory is:
@@ II.2 Doped bosonic Laughlin state
One may view $\varphi^{\dagger}$ as the creation operator of the anyon with charge $e/2$ in the standard effective theory of the Laughlin state. One may view $\varphi$ as the vortex of the composite boson. When the physical boson is at filling $n=\frac{1}{2}-x$, $\varphi$ is at density $n_{\varphi}=2x$. This is also in agreement with the doped CSL theory where the filling of the electron is n=12xn=1-2x so the Cooper pair density is $\frac{1}{2}-x$.
The next step is exactly the same as the doped CSL case. $\varphi$ feels an effective magnetic flux from $da$ and is at magnetic filling $\nu=-2$, so it can form a bosonic IQH phase with C=2C=-2. In the end the final theory is just:
@@ III Holon metal from doped SU(3)1 CSL
To describe the CSL, we use the slave boson construction $c_{i;a}=b_{i}f_{i;a}$ with $a=1,2,3$ as the flavor index. $f_{a}$ is put into a $C=1$ insulator with a $\frac{2\pi}{3}$ flux per unit cell. Similarly, the slave boson $b$ will feel a $2\pi/3$ flux per unit cell. Its dispersion will have three minima guaranteed by the projective translation symmetry. Next we dope the system so that the average density is $n_{b}=n_{f}=1-x$ per triangular unit cell. Similar to the SU(2) CSL case, $b$ and $f$ share an internal U(1) gauge field $a_{\mu}$. We assign the physical charge to the slave boson $b$. So in the end $b$ couples to $A_{c}-a$ and $f$ couples to $a$.
To keep $f$ in a $C=1$ Chern insulator, we need $\frac{\langle da\rangle}{2\pi}=-\frac{x}{3}$. Then the low energy theory is dominated by the slave boson $b$. Again $b$ feels an effective flux and its hole $\varphi\sim b^{\dagger}$ has three minima at momentum $\mathbf{k}_{1}=(0,0),\mathbf{k}_{2}=(\frac{2\pi}{3},0),\mathbf{k}_{3}=(\frac{4\pi}{3},0)$. We define the corresponding low energy field operator as $\varphi_{I}(t,x)$ with $I=1,2,3$. The physical translation symmetry acts as $T_{x}:\varphi_{I}\rightarrow e^{i\frac{2\pi}{3}I}\varphi_{I}$ and $T_{y}:\varphi_{1}\rightarrow\varphi_{2},\varphi_{2}\rightarrow\varphi_{3},\varphi_{3}\rightarrow\varphi_{1}$. The low-energy effective theory is:
@@
The translation symmetry acts as $T_{x}:\psi_{I}\rightarrow e^{i\frac{2\pi}{3}I}\psi_{I},\alpha_{\mu}\rightarrow\alpha_{\mu},\beta_{\mu}\rightarrow\beta_{\mu}$, $T_{y}:\psi_{1}\rightarrow\psi_{2},\psi_{2}\rightarrow\psi_{3},\psi_{3}\rightarrow\psi_{1},\alpha_{\mu}\rightarrow\beta_{\mu},\beta_{\mu}\rightarrow-\alpha_{\mu}-\beta_{\mu}$.
@@ III.1 Property of the holon metal
Here let us discuss the experimental property of the holon metal, with a focus on its transport signature. We hope to calculate the conductivity tensor $\sigma_{c}=\begin{pmatrix}\sigma_{xx}&\sigma_{xy}\\ -\sigma_{xy}&\sigma_{yy}\end{pmatrix}$. In the following, we assume that the translation symmetry is preserved. Thus, $\psi_{1},\psi_{2},\psi_{3}$ have the same conductivity tensor $\sigma_{\psi_{1}}=\sigma_{\psi_{2}}=\sigma_{\psi_{3}}=\sigma_{\psi}$. It is easy to find the response to $A_{c}$ as:
@@
where we used the Coulomb gauge with $A_{0}=0$ and $\mathbf{q}\cdot\mathbf{A}(\omega,\mathbf{q})=0$. There is no crossing term between $A_{c}$ and $\alpha,\beta$. $\Pi$ is the response from each $\psi_{I}$: $\Pi(\omega,\mathbf{q}=0)=\frac{e^{2}}{h}(-i\omega)\sigma_{\psi}=\frac{e^{2}}{h}(-i\omega)\begin{pmatrix}\sigma_{\psi;xx}&\sigma_{\psi;xy}\\ -\sigma_{\psi;-xy}&\sigma_{\psi;yy}\end{pmatrix}$. So in the end, we have a simple relationship: $\sigma_{c}=3\sigma_{\psi}$, provided that the translation symmetry is not broken. Therefore, the holon metal has a resistivity tensor which is very similar to a Fermi liquid. Because the time-reversal symmetry is broken, $\psi$ can still have a small non-universal Hall conductivity. But the final phase has a small Hall angle $\frac{\rho_{xy}}{\rho_{xx}}$ because $\sigma_{xx}$ is large in the clean system.
@@ III.2 Alternative derivation and wavefunction
To gain more insight into this exotic metallic phase, we also provide a different and perhaps simpler construction. In this phase, there are three pockets with Fermi surface size $A_{\mathrm{FS}}=\frac{x}{3}$ per Brillouin zone (BZ). We also find a gauge-invariant operator $C_{\mathrm{trion}}=\psi_{1}\psi_{2}\psi_{3}$ which is the physical trion operator. As discussed before, the physics is equivalent to doping a Laughlin state of the trion on top of a $C=3$ IQH phase. Now the trion is at density $n_{\mathrm{trion}}=\frac{1}{3}-\frac{x}{3}$, we simply do parton construction $C_{\mathrm{trion}}=f_{1}f_{2}f_{3}$ and then put each flavor $f_{I}$ in a C=1C=-1 Chern insulator plus a small hole pocket formed by $\psi_{I}\sim f_{I}^{\dagger}$ with size $x/3$. In this construction, there is a SU(3) internal gauge field, but let us assume it’s higgsed down to $U(1)\times U(1)$ by some terms. In the end we only have the U(1) gauge field $\alpha_{\mu}$ and $\beta_{\mu}$. $\psi_{1}$ couples to $A_{c}+\alpha_{\mu}$, $\psi_{2}$ couples to $A_{c}+\beta_{\mu}$ and $\psi_{3}$ couples to $A_{c}-\alpha_{\mu}-\beta_{\mu}$. The C=1C=-1 Chern insulator gives a background Chern-Simons term. Eventually, we reproduce the theory in Eq. 14 once we also include the $\frac{3}{4\pi}A_{c}dA_{c}$ term from the stacked $C=3$ IQH phase.